OpenStudy (anonymous):

The slope of a line is -3. Find the slope of a line that is perpendicular to this line. a.-3 b.-1/3 c.1/3

OpenStudy (anonymous):

1\3 perpendicular slopes have opposite intervals.

OpenStudy (dattatraya):

here given that slope = m = -3, we need to find the slope of line that is perpendicular ... the slope of perpendicular line is = -1/m , that is negative reciprocal of m. here m = -3, so slope = -1/m = -1/-3 = 1/3